OpenStudy (anonymous):

Prove: tan^2 θ cos^2 θ + cos^2 θ = 1. You must show all work.

hero (hero):

Hint: Factor out \(\cos^2\theta\)

OpenStudy (anonymous):

How do I do that?

hero (hero):

After factoring out \(\cos^2\theta\) you should have \(\cos^2 \theta (\tan^2 \theta + 1) = 1\)

hero (hero):

Next divide both sides by \(\cos^2\theta\)

OpenStudy (solomonzelman):

tan^2x=sin^2x/cos^2x cosines cancel and you get an identity.

hero (hero):

Or you can do that which would be better. @SolomonZelman just couldn't help himself.

OpenStudy (solomonzelman):

I don't really understand the last sentence of your comment :) -:(

OpenStudy (anonymous):

@SolomonZelman Can you help with one more

OpenStudy (solomonzelman):

When I "help" myself, posting the solution and answer, I get suspended. So why would I do that ?

OpenStudy (solomonzelman):

Gamer, Perhaps I can:) I'll try to flex my weak brains.

OpenStudy (anonymous):

I did a problem like that before with you but I seem to have forgot it @SolomonZelman

OpenStudy (solomonzelman):

attach not working -:(

hero (hero):

Actually, it would have been better to divide both sides by \(\cos^2 \theta\) because then you would have seen that \(\tan^2\theta + 1 = \sec^2 \theta\)

OpenStudy (anonymous):

@Hero Can you please help"?

hero (hero):

At that point you could have done this: \(\cos^2\theta(\sec^2\theta) = 1\)

hero (hero):

And since you know that \(\sec^2\theta = \dfrac{1}{\cos^2\theta}\) You could have written it this way for the next step: \(\cos^2\theta\left(\dfrac{1}{\cos^2\theta}\right) = 1\)

OpenStudy (anonymous):

@Hero Help with my other question

OpenStudy (anonymous):

Can you??

hero (hero):

Yes, but post it as a new question

OpenStudy (anonymous):

Sure but don't leave

OpenStudy (anonymous):

@Hero http://openstudy.com/study#/updates/538d3047e4b0d2ed7ed7c51e

hero (hero):

Try your best to post only one question at a time. Don't post two questions in the same discussion unless they are purposefully related to each other. Like for example Part A, and Part B, of Question 1